OpenStudy (anonymous):

Function Word Problem The basement of a large department store features discounted merchandise. Their policy is to reduce the previous month's price of the item by 8% each month for 6 months, and then give the unsold items to charity.

OpenStudy (anonymous):

i assume you're writing the rest of the question now, haha

OpenStudy (anonymous):

Let S6 be the sale price for the sixth month and P the original price. Express S6 as a function of P

OpenStudy (anonymous):

Yeah sorry @Omistone

OpenStudy (anonymous):

no worries, lemme write this out in my book so i can try to understand it a little clearer.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and then it says What is the final price of a $100 garment on sale for the sixth month? Round your answer to two decimal places.

OpenStudy (anonymous):

ok, so if we let P = the original price, and we know that the price decreases each month by 8% of the previous price, the price for any given month, n, can best be expressed as a geometric progression: \[S _{n} = P \times 0.92^{n-1} \] the reason we multiply by 0.92 is because 100% - 8% (for the discount) is 92% (0.92 in decimal form. so we simply let n = 6 \[S _{6} = P \times 0.904\]

OpenStudy (anonymous):

oh, right i didn't see your last post before i posted all that working out, but it should be easy enough for you to round it.

OpenStudy (anonymous):

so then , you just substitute in 100 for P in the equation and you will have your answer. I hope this helps, be sure to let me know if you don't understand anything.

OpenStudy (anonymous):

I think I understand it. Thank you

OpenStudy (anonymous):

S6=P×0.904 = 100 x 0.90 = 90 so the item would be $90 after 6 months

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

where does the minus 1 come from in \[0.92^n -1 \] @Omistone

OpenStudy (anonymous):

By what total percentage has the garment now been reduced from its original price? Round your answer to two decimal places.

OpenStudy (anonymous):

would the answer to the last one be 10% @Omistone